OpenStudy (anonymous):

find the value of the trig functions of theta from the info given cot theta =1/4, sin theta < 0 and tan theta=-4 , sin theta> 0

OpenStudy (anonymous):

what is the question?

OpenStudy (anonymous):

i just put it

OpenStudy (mertsj):



OpenStudy (mertsj):

\[\cos \theta =- \frac{1\sqrt{17}}{17}\] \[\sin \theta =-\frac{4\sqrt{17}}{17}\] \[\tan \theta=4\] \[\sec \theta=-\sqrt{17}\] \[\csc \theta=-\frac{\sqrt{17}}{4}\]

OpenStudy (anonymous):

okay thanks

OpenStudy (mertsj):

yw